La Silla Paranal Observatory News

The page contains important news and events of the La Silla Paranal Observatory. It also covers some important telescope and instruments news from the La Silla, Paranal and APEX sites.


General News

 

A Frequently Asked Questions page regarding ESO observations in the context of Covid-19 is available here http://eso.org/sci/observing/FAQ_COVID19.html

May 12, 2021 LPO update:

From May 17 to 28 we will increase on site-staffing at all Observatory sites, given the emerging (though yet somewhat fragile) improvement of the pandemic scenario in Chile.

This means in terms of systems to be operated

- La Silla both NTT and 3.6m (was only the NTT)

- APEX 16/24 hour operation (was 8/24)

- Paranal VLTI will be added, such that we foresee to do the VLTI-UT run starting on May 21 with all four UTs. Standalone UT operation during the shift May 17-28 will be limited to 2 UTs until May 21, and 3 UTs as of May 21.

 

May 5, 2021 Paranal update:

From May 7 to 17 we will continue with two UTs in operation. No VLTI-AT or survey telescope operations. The specific two UTs to be operated on a given night are decided based on operational restrictions and scientific priorities - no change between UTs during a given night. We expect to be able to execute all scheduled dVM observations on the UTs, and to have an overall preference for operations of UT3 & UT4.

 

April 27, 2021 Paranal update:

From April 26 to May 7, the two telescopes in operation are UT3 and UT4. N.B. the FORS2 visitor mode run originally scheduled in this period is rescheduled. dVMs on ESPRESSO are foreseen to be executed on UT3 or UT4.

 

April 19, 2021 LPO update:

From April 26 to May 7, Paranal will continue to operate only two UTs in a given night. No observations with VLTI or survey telescopes will take place. We plan to execute the dVMs on ESPRESSO and FORS2 scheduled during this period, on a best effort basis. 

La Silla remains operating only the NTT, and APEX in a 8/24hr scheme.

 

April 12, 2021 LPO update:

Due to the pandemic situation, operational activities at Paranal as of the shift April 16-26 will be further reduced. Paranal will operate two of the four UTs (baseline UT1 and UT4). VLTI-UT observations originally scheduled for April 20-28 will not take place. We plan to execute scheduled ESPRESSO dVMs on UT1 or UT4.

La Silla remains operating only the NTT, and APEX in a 8/24hr scheme.

 

April 5, 2021 Paranal update:

Due to the pandemic situation, Paranal Observatory operates only UT1, UT3 and UT4.

Out of operations are until further notice: VLTI observations with the ATs, the survey telescopes VST & VISTA, and UT2. An intervention to fix the passive support of UT2 is tentatively scheduled for early May.

It is foreseen to operate VLTI with the UTs from April 20-28. 

March 26, 2021 Paranal update:

The operation of UT2 continues to be affected by failure of its passive support. For VLTI-UT observations, UT2 operability could be recovered via a workaround. Therefore, the current operations on Paranal cover the ongoing VLTI-UT run (until April 2).

March 24, 2021 Paranal update:

Due to the degradation of the sanitary situation, Paranal will not operate the survey telescopes (VISTA, VST) until further notice. In addition, a major technical problem affects UT2, and therefore also prevents operations with UT2 and the VLTI; this problem will not be fixed before April 6th.

March 24, 2021 La Silla update:

Due to degradation of the sanitary situation in the Region of Coquimbo, La Silla will only operate the NTT from the night of March 24th until at least April 6th.

 

March 12, 2021 Paranal update:

Update on KMOS: arm 22 had failed and it is now stuck within the field of view. To prevent collisions with other arms, KMOS is currently operating in a degraded mode with 18 arms available. USD will provide affected users with instructions to revise their observations.

 

March 4, 2021 Paranal update:

HAWK-I recommissioning will not happen before April. 

At the time of writing, KMOS@UT1 is out of operations due to a problem with one specific arm that affects its operational readiness. Efforts are ongoing to address the problem at the technical level and the level of observation preparation.

 

February 17, 2021 Paranal update:

Unfortunately, further contamination in the HAWK-I entrance window was detected, and also oil condensation on one of the filters. HAWK-I is not going to be recommissioned these days and needs to be warmed up again. Recommissioning is suspended for several weeks more, at minimum.

 

February 10, 2021 Paranal update:

Our second on-sky recommissioning of HAWK-I, after fixing the entrance window contamination, is planned to take place around February 23. 

 

January 28, 2021 Paranal update:

After an initial test on sky on January 17, significant contamination was found on the entrance window of HAWK-I. Recommissioning of HAWK-I is suspended until later in Q1.

The first week of on-sky commissioning of CRIRES+ was completed on January 26.

The operational systems in Paranal are:

FORS2 & KMOS@UT1, UVES & FLAMES @UT2, SPHERE & X-Shooter@UT3, MUSE@UT4 with AOF, ESPRESSO@ICCF, OmegaCAM@VST, VIRCAM@VISTA, Gravity/MATISSE/PIONIER @ VLTI.

 

January 16, 2021 Paranal update:

KMOS@UT1 and FLAMES@UT2 are back in regular operations.

 

January 10, 2021 Paranal update:

VLTI-UT is back in regular science operations since January 5.

KMOS@UT1 and FLAMES@UT2 are under active recommissioning and expected to become available soon.

 

December 17, 2020 Paranal update:

VIRCAM@VISTA started science operations on December 15th, bringing back all Paranal telescopes. The operational systems on Paranal are:

FORS2@UT1, UVES@UT2, SPHERE & X-Shooter@UT3, MUSE@UT4 with AOF, ESPRESSO@ICCF, OmegaCAM@VST, VIRCAM@VISTA, Gravity/MATISSE/PIONIER @ VLTI.

In January we plan to add VLTI-UT capabilities, KMOS@UT1, FLAMES@UT2, and HAWK-I at UT4.

 

December 14, 2020 Paranal update: MATISSE back

On the night of December 13-14, MATISSE successfully observed Betelgeuse in a visitor mode run. Coincidentally, this is 100 years to the day after its diameter was measured for the first time using a beam interferometer invented by Michelson.

 

December 12, 2020 Paranal Update: ESPRESSO, MATISSE, next systems 

ESPRESSO has completed temperature stabilisation and is now fully operational also for highest accuracy RV monitoring programmes.

MATISSE recommissioning is completed in terms of instrument and system performances. Final adjustments regarding the end-to-end data flow are being made and we expect MATISSE ready for science observations by tomorrow Sunday, December 13.

First on-sky tests with VISTA & VIRCAM are planned for December 14/15.

On-sky recommissioning of FLAMES was planned until December 15, but will likely be postponed by up to a few weeks.

 

December 8, 2020 Paranal Update: ESPRESSO back for non-RV programmes    

ESPRESSO has restarted science observations for non-RV monitoring programmes. Full stabilisation to allow most accurate RV monitoring is still ongoing. MATISSE is undergoing on sky commissioning and foreseen to be science ready in the next nights.

 

December 2, 2020 Paranal Update: Gravity@VLTI back, ESPRESSO on track

Gravity is back in regular science operations since November 30. ESPRESSO is on track to restart science observations in the next nights. MATISSE needs a few more nights to start on-sky recommissioning.

UT3 and UT4 have been out of service for 3 and 4 nights, respectively. UT3 due to a problem with the hydraulic bearing system, UT4 due to a problem with an optical element for image analysis. Both telescopes are back on sky in nominal performance since the night December 1/2. 

 

November 27, 2020 Update Paranal: PIONIER@VLTI back

Since November 26, observations with PIONIER@VLTI have resumed. Gravity is under active recommissioning.

 

November 23, 2020: Update Paranal: VST back, further plans

Since November 17, observations with OmegaCAM@VST have resumed.

We are currently bringing the VLTI facility and instruments back to operations. First instrument on sky is expected to be PIONIER in the next days, followed by Gravity/MATISSE next week.

Early December we plan to have ESPRESSO back. Until middle of December we plan to add VIRCAM@VISTA and FLAMES@UT2 to the systems in operations. 

For January we foresee a sequence of VLTI-UT, KMOS (both early January), HAWK-I (later January).

 

November 13, 2020: Update Paranal: SPHERE and MUSE-AOF back 

Since November 12/13, SPHERE and MUSE-AOF (WFM & NFM) are back in regular science operations.

Assuming that further restart activities go smoothly and the weather collaborates, the next systems expected to become ready until end of November are OmegaCAM and the VLTI instruments. 

 

November 11, 2020: Update Paranal on next systems to go on sky

Current systems in operations: FORS2, UVES, X-Shooter, MUSE without AO.

We expect the following systems to become ready for science operations during the next week: MUSE-AOF, SPHERE, OmegaCAM@VST. We will shortly provide an update on the sequence to be expected after this, with the goal to have most systems in operations by the end of the year.

 

November 11, 2020: Update on La Silla Status

Observing with NTT/EFOSC2 since October 19

Observing with 3.6m/HARPS since November 9.

NTT/SOFI foreseen before end of November.

 

October 29, 2020: Update Paranal & La Silla: MUSE and EFOSC back in operations

On the night of October 26, first scientific data were taken in Paranal with MUSE @ UT4, which is now back in regular operations. At present, only no-AO observations are supported. We plan to add AOF capabilities in the next weeks.

On the night of October 28, first scientific data were taken at La Silla with EFOSC @ NTT, in designated visitor mode.

 

October 15, 2020: Update Paranal: X-Shooter back in operations

On the night of October 12, first scientific data were taken with X-Shooter which is now back in regular operations.

 

October 1, 2020: Update Paranal: next instruments to become available

The current plans foresee to be ready with X-Shooter@UT3 by October 21, and MUSE@UT4 _without adaptive optics_ by November 4.

 

September 18, 2020: Update Paranal: science data being taken again with both FORS2@UT1 and UVES@UT2

In the evening of September 17, UT1 went back on sky for the first time since March. After testing the functioning and performance of both telescope and instrument during the night, first science data with FORS2 were taken at the end of the night. 

Service mode observations with both UVES and FORS2 are thus running again.

 

September 15, 2020: ramp-up to minimal science operations has started on APEX and Paranal. First scientific data on Paranal has been taken.

On September 9, the ramp-up towards minimal operations at APEX and Paranal was started, with the first shifts arriving to both sites.

On September 13, first scientific data were taken for UVES@UT2. FORS2@UT1 is being prepared for scientific operations and is expected to become available in the next days.

At APEX the ramp-up is also going according to plan and the re-start of scientific observations is imminent.

 

August 28, 2020: Minimal science operations to be resumed at APEX, Paranal and La Silla, see also https://eso.org/sci/publications/announcements/sciann17325.html

Following a detailed assessment of the current situation and the logistic, technical, and safety constraints, as well as the capacity to guarantee the compliance with all measures needed to protect the health of its staff at all times, ESO has concluded that it will be able to re-establish a minimum level of science operations at the observatory sites in the coming weeks.

Under the minimal science operations scheme, the following activities will be resumed:

  • Paranal Observatory: science operations at the VLT with FORS2 at UT1 and UVES at UT2, in Service Mode and Delegated Visitor Mode.
  • La Silla Observatory: science operations at the NTT with EFOSC2, in Delegated Visitor Mode. The possibility of adding HARPS at the 3.6m telescope is still under consideration.
  • APEX: science operations with SEPIA at frequencies below 500 GHz. The possibility of adding nFLASH or PI230 is still under consideration.

The expected dates of re-starting operations under this scheme are 23 September at Paranal and APEX. It is not yet possible to provide an estimated restart date for La Silla, which would take place in any case after Paranal. The observations to be performed will be selected from the backlog accumulated during the past months. Please note that the Director's Discretionary Time channel remains closed at this time.

ESO will continue to closely follow all the developments influencing its ability to proceed with the planned resumption of minimal science operations and eventual further expansion to include more instruments and telescopes, and it will revert back to the present safe mode if health and safety considerations so dictate.

Further inquiries on this matter can be addressed to the ESO User Support Department.

 

August 10, 2020: UPDATE due to Corona virus:

All ESO observatories continue in safe state, no science operations are taking place. 

The sixth review to assess whether the conditions in Chile allow to initiate the ramp-up plan took place last week. It has resulted in the decision to not initiate the ramp-up at this moment. 

As a consequence, visitor mode runs until end of September can be cancelled. Users will be informed individually by ESO.

We will update this page bi-weekly with the outcome of the regular reviews.

 

July 28, 2020: UPDATE due to Corona virus:

All ESO observatories continue in safe state, no science operations are taking place. 

The fifith review to assess whether the conditions in Chile allow to initiate the ramp-up plan took place last week. It has resulted in the decision to not initiate the ramp-up at this moment. 

As a consequence, visitor mode runs until September 10 can be cancelled. Users will be informed individually by ESO.

We will update this page bi-weekly with the outcome of the regular reviews.

 

July 13, 2020: UPDATE due to Corona virus:

All ESO observatories continue in safe state, no science operations are taking place. 

The fourth review to assess whether the conditions in Chile allow to initiate the ramp-up plan took place last week. It has resulted in the decision to not initiate the ramp-up at this moment. 

As a consequence, visitor mode runs until end of August can be cancelled. Users will be informed individually by ESO.

We will update this page bi-weekly with the outcome of the regular reviews.

 

June 29, 2020: UPDATE due to Corona virus:

All ESO observatories continue in safe state, no science operations are taking place. 

The third review to assess whether the conditions in Chile allow to initiate the ramp-up plan took place last week. It has resulted in the decision to not initiate the ramp-up at this moment. 

As a consequence, visitor mode runs until middle of August are cancelled. Users will be informed individually by ESO.

We will update this page bi-weekly with the outcome of the regular reviews.

 

June 17, 2020: UPDATE due to Corona virus:

All ESO observatories continue in safe state, no science operations are taking place. 

The second review to assess whether the conditions in Chile allow to initiate the ramp-up plan took place last week. It has resulted in the decision to not initiate the ramp-up at this moment. 

As a consequence, visitor mode runs until end of July are cancelled. Users will be informed individually by ESO.

We will update this page bi-weekly with the outcome of the regular reviews.

 

June 3, 2020: UPDATE due to Corona virus:

All ESO observatories continue in safe state, no science operations are taking place. 

The first review to assess whether the conditions in Chile allow to initiate the ramp-up plan took place last week. It has resulted in the decision to not initiate the ramp-up at this moment. 

As a consequence, visitor mode runs until middle of July are cancelled. Users will be informed individually by ESO.

We will update this page bi-weekly with the outcome of the regular reviews.

 

May 22, 2020: UPDATE due to Corona virus:

All ESO observatories continue in safe state, no science operations are taking place. All visitor mode runs until end of June have been cancelled.

LPO has developed a ramp-up plan to transition the Observatories from Safe State to Restricted Operations mode, with the ability to perform basic science operations with limited staffing on site.

As of the week May 25, bi-weekly reviews will be held to assess whether the conditions in Chile allow to initiate the ramp-up plan. Once the ramp-up plan is initiated, we estimate a minimum of three weeks until first scientific data can be taken again, and about two more weeks until 'steady state' restricted operations would be reached.

An update will be posted here as soon as the ramp-up plan is triggered.

 

May 7, 2020: UPDATE due to Corona virus:

All ESO observatories continue in safe state, no science operations are taking place. In addition to April and May, also all visitor mode runs for the first two weeks of June have been cancelled.

An update will be posted here as soon as the ramp-up plan to restart science observations is triggered. For the time being, no date has been determined for this trigger.

 

April 22, 2020: UPDATE due to Corona virus:

All ESO observatories continue in safe state, no science operations are taking place. All visitor mode runs for April and May have been cancelled.

An update will be posted here as soon as the Observatories restart their activities.

 

March 24, 2020: UPDATE due to Corona virus: 

As anticipated in the news item from March 19, all ESO observatories have stopped science operations by now.

  • Paranal collected its last photons on the night of March 23/24, with FORS2, UVES, SPHERE, MUSE and PIONIER as last instruments online.
    • By March 25, Paranal will have moved to a Safe State with all telescopes and instruments switched off, only essential systems for monitoring and emergencies online, and a minimal staffing on site of about 20 people (compared to 150 during operations).
  • La Silla had its last night of science operations on March 22/23. All ESO telescopes (3.6m, NTT, 2.2m) and instruments have been switched off. Also hosted systems have been ramped down. A minimal staffing on site of 5 people will remain.
  • APEX went into 'survival'/mininmal mode by March 21. Telescope and instruments have been shut down and secured. At the high site, only essential systems are running with an estimated power autonomy of eight more weeks. Sequitor base camp has been secured and has supplies for the next 6-8 weeks. 
    • Three staff will remain at the Sequitor base camp during survival mode. 

 

March 19, 2020: UPDATE due to Corona virus: 

To minimise the risk of infection the presence of staff has been significantly reduced at the observatories over the past week leading to a number of restrictions to the ongoing science operation.

Over the coming weekend, the staff at the observatories will be further reduced to the minimal teams to ensure the safety of the facilities and the remaining people on site. 

As a consequence science operation will cease at all sites:  

·         On Paranal, the last night of science operation is foreseen for 23/24 March 2020. Selected instruments and telescopes will cease science operation before.

·         On La Silla, the NTT and 3.6-metre telescopes will stop science operation after the night of 22/23 March 2020.

·         At APEX, science operation will not restart after the planned shutdown period that lasted from December until now.

 

March 16, 2020: Reduced Operations due to Corona Virus.

In response to rapidly evolving COVID-19 pandemic around the world ESO has been following the measures adopted by governments in Germany and Chile and closely observing the situation in other countries. Our highest priority remains the safety and wellbeing of all staff and visitors at ESO’s sites.

With the latest developments, which includes travel restrictions to and from Chile and Europe decided on government levels, ESO has canceled all international travels to Chile until at least 19 April 2020.

To minimise the risk of infection the presence of staff is significantly reduced at the observatories. 

As a consequence, the operations on the sites gradually move to a restricted mode as of March 16:

  • On Paranal, Service Mode and designated Visitor Mode observations on UTs and the VLTI will continue to be supported, on a “best effort” basis including remote support by staff astronomers from their home.
    • Operation of the survey telescopes VST and VISTA have the lowest priority.
  • On La Silla Visitor Mode runs with observers coming from overseas are cancelled until 19 April, and ESO will evaluate on a case-by-case basis if any of the scheduled observing runs could still be executed remotely. 

  • APEX operations are coordinated with scientists from all partners to keep remote observing with ESO staff on site to the extent possible. 

The above situation may change quickly. Updates of the status of the observatory will be posted at this page. For urgent questions please contact usd-help@eso.org (service mode) and paranal@eso.org (visitor mode). For La Silla please contact lasilla@eso.org. For APEX please contact apex-garching@eso.org.


Paranal News

Telescope News

2019

  • UT3
    • January: SINFONI installed on cassegrain focus
  • UT4
    • January: the telescope loses SINFONI waiting for the NEAR experiment

2018

  • UT3
    • December: VISIR moved from cassegrain focus to NIH for NEAR upgrades
  • UT4
    • May: Altitude motor repair
    • March/April: DSM binary chopping test

2017

  • All UTs
    •  November: TCS VLT2017 upgrade
  • UT2
    • May: Enclosure Rotation drives upgrade
  • UT3
    • December: Enclosure PLC upgrade
    • October: Altitude Drives upgrade
    • May/June: Espresso Coudé train installation
    • April: M1/M3 recoating
  • UT4
    • November-December: LGSF maintenance and LPC swap
    • October-December: GRAAL comissioning runs
    • May-September: GALACSI WFM commissioning
    • August: M1/M3 re-coating
    • May: Altitude Drives Upgrade
    • April
      • realignment of Nasmyth B assembly of Adapter/Rotator, its support structure and GALACSI. 5 tons of equipment and steel moved of ~1.5 mm without dismounting from the telescope fork structure (TTR 99.0045)
      • completion of first phase of AOF-GALACSI commissioning
    • March: UT4 Naysmith B Drive Motor: removal of magnet protection plates and out-of-service cooling lines
    • February
      • UT4 altitude motor segment replacement (TTR-098.0077)
      • completion of first phase of AOF-GRAAL commissioning

2016

  • UT2
    • 7-12 December: M1/M3 recoating
  • UT4
    • 2 December: return to Science operations after DSm integration
    • 24 October: first light with DSM after its integration

2013

  • 20th November: recoating of M1, M3 of UT2 completed
  • 11th September: recoating of M1, M3 of UT4 completed
  • 1st April: The VST is currently out of service between March 29, 2013, and April 2, 2013 due to a multiple damage of cables and cooling lines after an incident caused by the failure of a limit switch. All damaged parts were replaced or repaired.
  • 26th February: The Laser Guide Star system at UT4 has recently been successfully upgraded with a fibre Raman laser (AOF prototype) delivering improved on-sky performances as well as enhanced operational stability and maintenance. Full characterization of the upgraded system will soon be available. This intervention will result in a dramatic increased in the availability of the LGS system for science operations.

 

Instrument News

2019

  • June 25th: SINFONI has been warmed up for its decommissioning. SPIFFI will come back next year after upgrade and integration within ERIS.
  • May: KMOS: During an intervention, arms #4 and 8 have been replaced. However, arm#14 showed a hard point and was left in simulation in park position.
  • May 7-12: X-shooter is unavailable due to a Cassegrain intervention.
  • April 19: KMOS arm #8 is stuck and deployed in the field-of-view at its maximum linear extent.  

2018

  • November 27 to December 7: X-shooter is not available due to maintenance.
  • September 23: An earth tremor caused a shift in the direction perpendicular to the dispersion direction of ~4 pixels in the HR grating only.  An intervention on 13-Oct-2018 returned the grating to its original position.
  • October 1: All observations should now be prepared with p2; visiting astronomers in the control room can also use vOT. See https://www.eso.org/sci/observing/phase2/p2intro.html.
  • October 1: FORS2: the high time resolution modes (imaging, HIT-I; spectroscopy, HIT-S; and multi-object spectroscopy, HIT-MS) are no longer offered.
  • October 1: KMOS arm #4 has started to fail in all its movements since September 26; it was placed in simulation.
  • August: KMOS: Failure of a cold head leads to an emergency intervention to replace it. Following the intervention, the performances of spectrograph #2 are degraded with astigmatism along the spatial direction  and spectral direction with and degraded resolution of ~10%.
  • July: NACO: Due to unrecoverable problems, the visible wavefront sensor was decommissioned.
  • May: KMOS intervention to maintain the cold heads and fix/replace the underperforming arms (namely arm #1, #3 and #17). Arm 1 and 3 have been replaced and arm 17 fixed
  • April 12 - May 24th: NACO: polarimetry has not been available  due to an issue with the retarder.
  • March: NACO had several issues with the CONICA detector.
  • Janury 7: KMOS arm #17 also shows a rotator-angle dependent flux loss since December 2018.

2017

  • November 14: KMOS arm #1 shows flux loss of up to 40%. In addition, the flux loss depends on the rotator angle.
  • October: UVES: Second and final part of the intervention on the blue camera lenses, to recover the resolution of the UVES blue arm, is now completed. The resolving power in the UVES blue arm is back to its early-days values, with a most-recent median measurement of R~93,000 (for the 0.4-arcsec slit).
  • April: NACO: the CONICA detector is very unstable.
  • February: UVES: An intervention to improve the resolution of the blue arm, which has shown to degrade slowly through the years due to aging of the blue camera, took place. The resolving power measured after the intervention is ~85,000 for the 0.4-arcsec slit. The RMS of the residuals of the wavelength calibration improved by a factor of two.
  • January 9: KMOS: Arm 3 is still in simulation and will remain in this state as soon as the instrument can be warmed up and the arm fixed/replaced. Even though the astrometry is indeed re-calibrated, Arm 3 is still showing vignetting on sky with such vignetting being dependent on the arm and rotator position.

2016

  • December 20: X-shooter: After X-shooter was remounted following the last mirror recoating, the readout noise is much higher in the VIS arm. All observations since 2016-12-20 are affected. The readout noise increased by 20-70% depending on the readout mode used, see http://www.eso.org/observing/dfo/quality/XSHOOTER/reports/HEALTH/trend_report_BIAS_VIS_ron_raw_HC.html. Unfortunately, the cause has not yet been identified.
  • December 19: UT4 instruments. Following the installation of the deformable secondary mirror, all instruments on UT4 have now been recommissioned and are back in operation.
  • December 5: NACO: The first part of an effort to re-align NACO has been completed. The field stop was re-aligned and the number of AO modes significantly reduced. Performance using the VIS wavefront sensort is nominal; additional work will be done in January 2017 for the IR WFS.
  • November 25: VIMOS: The IFU intervention was completed and we have been observing different specphots to adjust the center of the IFU. The pointing accuracy is now within operational limits given the blind IFU pointing (3 arcsec).
  • November 12: VIMOS: The IFU unit is currently inoperative due to a problem with the IFU shutter. An intervention will take place shortly, most likely before Nov 17.
  • October 9: All UT4 instruments are unavailable due to the installation of the Deformable Secondary Mirrors. Operation will restart in December.
  • October 1: GRAVITY enters operation with the auxiliary telescope array in spectro-imaging mode for the medium and high spectral resolution settings. Gravity can be used in both single and dual field modes (i.e. both on-axis and off-axis fringe-tracking).
  • September 13: SINFONI: The MACAO shutter for the AO unit was replaced successfully.
  • September 5: KMOS: Arm 3 is experiencing again issues: bad precision in positioning as before plus vignetting. These issues are under investigation and Arm 3 is set to simulation.
  • July 11: FLAMES: A new version of FPOSS is available.
  • July 8: KMOS: arm #3 shows problems. It had been set to simulation for a short period (25th July to 4th August)  but is now working again, albeit with a higher uncertainty on the positioning (about 3 pixels). For that reason, the arm to take the telluric standard for spectrograph 1, was changed from #3 to #4.
  • May 18-22: OmegaCAM was out of operations for maintenance.
  • April 2: UVES: An intervention to restore the full performance of the blue arm took place at the period change over to P97. The blue-arm camera is now back to its condition pre-earthquake, i.e., the blue arm delivers a resolving power of ~75,000 for a slit width of 0.4 arcsec.
  • March 31: SINFONI has been upgraded. See the News page for further information.
  • Feburary 27: A new Giraffe readout mode will be availble in Visitor Mode only from Period 97. The mode is a high gain, low noise, slow readout mode that results in higher S/N (about a factor 2) only for faint targets, i.e. at the limit of the instrument sensitivity. The readout time of this slow mode is about 180s, whereas in the standard readout mode this is 45s. The ETC already includes the new readout mode in the Instrument Setup section. More details will be available in the User and Template manuals.
  • February 21: The PSF was fully recovered on February 18th. Image quality and contrat performances are now back to nominal.
  • Feburary 5: VIRCAM: The response in the upper half of chip #16 has decreased after the intervention in Nov 2015. The effect is more pronounced in the Z band, and to a lesser extent in Y, J and the narrowband filters. Before the intervention, the fractional response in the upper half was ~50%-70% in Z. After the intervention it has decreased to ~40%. This issue is under investigation.
  • February 4: SPHERE:  Since January 19th, SPHERE suffers from a degraded image quality (a factor ~0.9 to the H-band Strehl ratio and a factor ~0.5 to the R-band Strehl ratio). The PSF core has a "cross" like shape". The team is working on finding as solution. SPHERE is kept in operations.

 


La Silla News

The section provides the latest telescope and instruments news from the La Silla Site.

More news can be found in the news section of theLa Silla Science Operations page.

  • March 14, 2010: Updated service observing information Service Observing
  • November 20, 2009: A new technique using near-infrared images, obtained with s 3.58-metre New Technology Telescope (NTT)ESO Watching a Cannibal Galaxy Dine
  • April 6, 2009: La Silla Astroclimatology information (weather, seeing, extinction) was moved to a new location and can now be found on the mean navigation menu of the La Silla Facilities pages. Additional La Silla astroclimatology information will be added here.
  • February 11, 2009: From the second week of July, 2009, observations will be carried out from a new control room located in theadministration building.
  • February 11, 2009: A new page describing the way that operations will run after P83 (ie from October 2009) is available here.
  • April 22, 2008: A document that explains how to interface visitor instruments with focal stations of La Silla telescopes is available (2.8 MB PDF format).
  • March 27, 2007: The dome of the 3.6m has been fully restored, and moves with nominal speed.
  • October 05, 2006: The dome of the 3.6m enters an extended period of corrective maintenance. Science operations at the 3.6m is temporarily suspended until further notice.
  • May 17, 2006: A Trio of Neptunes  discovered with HARPS, see ESO PR 18/06
  • January 26, 2006: A 5.5M_earth planet discovered with the 1.54m Danish telescope.